OpenStudy (jpsmarinho):

How I can understand this: Lim(x*sin(1/x)) when x goes to 0 equals to 0 ? Thanks!

OpenStudy (anonymous):

sine fluctuates very quickly, because as \(x\to 0\) you have \(\frac{1}{x}\to \infty\) but sine is stuck between -1 and 1, so \(x\sin(\frac{1}{x})\) is stuck between \(-x\) and \(x\)

OpenStudy (anonymous):

and as \(x\to0\) clearly \(-x\to 0\) as well, so the whole thing goes to zero here is a picture www.wolframalpha.com/input/?i=xsin(1%2Fx) you can sort of see that the function \(x\sin(\frac{1}{x})\) is bounded by the two lines \(y=x\) and \(y=-x\)

OpenStudy (jpsmarinho):

This is the confrontation theorem, right?

OpenStudy (anonymous):

i have never heard of the that theorem, but maybe. sometimes i have heard this called the "squeeze" theorem, but for all i know they are the same

OpenStudy (jpsmarinho):

I'm Brazilian, here it's called confrontation, but I've heard squeeze. Thanks!!